OpenStudy (anonymous):

Solve for x: x^2 < |x|

OpenStudy (anonymous):

There is a formula for this, but I dont do it that way. you can set that equal to x^2=x and x^2=-x. solve for x in both and test the values you find, plus the regions in between and see if it holds true in the original problem. then you can write it in interval notation. Let me know if none of that made sense with what you've been doing in class

OpenStudy (richyw):

\[-1<x<1\]

OpenStudy (anonymous):

I did the method where you solve x^2=x and x^2=-x, and i had the values: 0, -1 and 1 . I want to know how it is -1<x<1 why x cannot equal 0.

OpenStudy (anonymous):

x cannot equal 0,1,-1 based on the original equation and there is a hole at that point. As for why it doesnt work conceptually beyond the limitation the equation gives....i dont know, thats a question for your instructor.

OpenStudy (anonymous):

so its (-1,0) U (0,1)

OpenStudy (richyw):

right. missed that.